Hello, the equation I have is a word problem which I have a really hard time with.
Below is the problem:
The amount of money Jim spends on fishingdepends on the average temp during any mouth. It varies according to the function m=5t^3/5+45. If the temp in July was 87.45 degree how much did Jim spend?

Answer by checkley77(12569)

m=5t^3/5+45
m=5*87.45^3/5+45
m=5*14.624+45
m=73.12+45
m=$118.12 ans.